please help asap (please make answer clear). please provide all missing answers to the blank spaces
Find and interpret f(2) and f(2). f(x) = - f(2) = Preview f(2) = Preview So at x = 2, f(x) is Select an answer at a(n)
Answer #1

Truyo. Date: f(x)= -3 SC 7(x) = 3 Hence t(2) = 3 +168.2 = - 3 X2.. X3 -6 x 3 +1(x) = +2) = - 43 t (2) = -6 -6 - 3 23 4 |